OpenStudy (anonymous):

using quadratic formula solve: x^2+7x=-3

OpenStudy (anonymous):

start with \[x^2+7x+3=0\] and use the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]with \(x=1,b=7,c=3\)

OpenStudy (anonymous):

im stuck on this part \[x= -7\pm \frac{ \sqrt{49}+12 }{2}\]

OpenStudy (anonymous):

@satellite73

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

what do i do next?

OpenStudy (jdoe0001):

-4ac => -4 * 1 * 3 \(\ne +12\)

OpenStudy (jdoe0001):

\(\bf \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\implies x= \cfrac{ - 7 \pm \sqrt { 7^2 -4(1)(3)}}{2(1)}\)

OpenStudy (jdoe0001):

so... what would you get inside the radical?

OpenStudy (jdoe0001):

hmm anyhow \(\bf \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\implies x= \cfrac{ - 7 \pm \sqrt { 7^2 -4(1)(3)}}{2(1)}\\ \quad \\ x= \cfrac{ - 7 \pm \sqrt { 49 -12}}{2}\implies x= \cfrac{ - 7 \pm \sqrt { 37}}{2}\implies \begin{cases} \cfrac{ - 7 + \sqrt { 37}}{2}\\ \quad \\ \bf \cfrac{ - 7 - \sqrt { 37}}{2} \end{cases}\)

OpenStudy (anonymous):

you would get \[-7\pm \sqrt{49}+12\] right?

OpenStudy (jdoe0001):

\(\bf -4(1)(3)\implies -12\)

OpenStudy (anonymous):

thank you!!! i see what i did wrong

OpenStudy (jdoe0001):

yw